OpenStudy (anonymous):

Simplify -3x square root sign 20x^2

OpenStudy (anonymous):

@Diamondsa Is this your problem? \[-3x \sqrt{sign(20x^2)}\]

OpenStudy (cornzyblack):

You mean \[-3x \sqrt{(20x ^{2}}\] that is, \[-3x \sqrt{4 * 5 * x ^{2}}\] then what we have is \[(-3x * 2 * x) \sqrt{5}\] which is\[-6x ^{2} \sqrt{5}\]

OpenStudy (anonymous):

I agree with CornzyBlack.

OpenStudy (cornzyblack):

thankYou :)